OpenStudy (anonymous):

Given = xy -x^2y, x = s + t, y = s - t...Find Zs and Zt

OpenStudy (anonymous):

Use chain rule\[\large z_s = z_x\cdot x_s \]

OpenStudy (anonymous):

\[ z_x = y-2xy \]\[ x_s = 1 \]\[ z_s = z_xx_s = y-2xy = (s-t)-2(s+t)(s-t) = s-t-2(s^2-t^2) \]

OpenStudy (anonymous):

Whoops... actually I should have put \[ z_s=z_xx_s+z_yy_s \]

OpenStudy (anonymous):

\[ z_y = x-x^2 \]

OpenStudy (anonymous):

\[ y_s = 1 \]

OpenStudy (anonymous):

so you have \[ z_s = (y+2xy) + (x-x^2) \]

OpenStudy (kc_kennylau):

Please add a \Large command before every LaTeX line... the subscripts cannot be clearly seen without zooming the browser...